OpenStudy (anonymous):

is there any equivalent number of sine so that i can solve sine 70 without using any calculator? tnx

OpenStudy (shubhamsrg):

sin 3x = 3sin x - 4 sin^3x so sin 210 = 3 sin 70 - 4sin^3 70 =>-sin 30 = 3 sin70 - 4sin^3 70 -> 8sin^3 70 - 6sin 70 -1 =0 let sin70 =x => 8x^3 - 6x -1 =0 that'd now require some higher mathematical skill to solve it..but you dont have to use much of calculator ;)

OpenStudy (anonymous):

i'm sorry, but how did you get sin 3x = 3sin x - 4 sin^3x ? thank you :)

OpenStudy (shubhamsrg):

thats a universal formulla .. you may derive it through many paths here's one note that sin(a+b) = sina cosb + cosa sinb will this help ?

OpenStudy (anonymous):

oh i see, thank you :) it help me so much :))

OpenStudy (shubhamsrg):

glad to help ^_^